Circular No. 7531

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ey IN IC 1481 R. Chornock, W. D. Li and A. V. Filippenko, University of California,

Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Nov. 11.2 (mag ∼ 16.2), 20.2 (mag ∼ 15.4) and 28.2 UT (mag ∼ 15.6). The new object is located at α = 23^{h}19^{m}25^{s}.20, δ = +5^{o}54′21′′.1 (equinox 2000.0), which is 0′′.6 east and 1′′.2 south of the nucleus of IC 1481. A KAIT image of the same field on Nov. 6.3 showed nothing at the position of the supernova (limiting mag ∼ 18).

Filippenko and Chornock report that CCD spectra (range 330–1000 nm) obtained on Nov. 29 with the Shane 3-m reflector at Lick Observatory show that the object is a normal type Ia supernova, 1 ± 2 days past maximum brightness (using the spectral-feature aging method of Riess et al. 1997, A.J. 114, 722).

SUPERNOVA 2000el IN NGC 7290 Chornock, Li and Filippenko also report the independent discovery by

LOTOSS of SN 2000el (IAUC 7523). Available photometry (all unfiltered), obtained after subtracting a template image obtained on 1999 Nov. 25: 2000 Oct. 18.2 UT, [19.0; 23.1, 17.2; 31.1, 16.7; Nov. 8.2, 17.0; 11.2, 17.0; 27.1, 16.8. Filippenko and Chornock report that Nov. 29 spectra obtained as above show that the object is a normal type II supernova. Hα has a classic P-Cyg profile, and there are several prominent Fe II absorption lines.

KS 1947+300 = GRO J1948+32 J. Swank, Goddard Space Flight Center; and E. Morgan, Massachu-

setts Institute of Technology, report that the transient x-ray source KS 1947+300 (IAUC 7523) is a pulsar with a barycentric period on Nov. 21 of 18.7579(5) s. This period is consistent with the object’s being identical with the pulsar GRO J1948+32 (Chakrabarty et al. 1995, Ap.J. 446, 826), having slowed down from the 18.70 s of an outburst in 1994 at an average rate of 8 ms/yr. The flux was 4.5 × 10^{−10} erg cm^{−2} s^{−1} above 2 keV, with a cutoff power-law spectrum (photon index ∼ 0.6, with a 10-keV e-folding energy above 6 keV) and a small column density (< 10^{21} cm^{−2}). A sequence of RXTE observations was undertaken, with the source last seen on Nov. 28 at a flux diminished by a factor of four.

2000 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden